# Quantum anomalous Hall phases in gated rhombohedral graphene

### v1

@@ 2 Effective model and classification
for $0\leq k\leq\lfloor\frac{m}{2}-1\rfloor$ and $\delta_{\lfloor\frac{m}{2}\rfloor}=\infty$.
Setting $j=k+\lceil\frac{m}{2}\rceil$ and rearranging (11) in terms of $\delta_{k}$ we deduce that phase transitions happen at critical points of potential $u$, assuming $\gamma$ remains constant:
$u_{k}=\gamma\frac{m-1}{\sqrt{\delta_{k}^{2}-\frac{1}{4}}}$ (13)
for $1\leq k\leq\lfloor\frac{m}{2}-1\rfloor$. We observe the existence of only one phase when $m=2$ or $m=3$. Assuming $m\geq 4$, the smallest value of $u$ at which we may observe a transition for a fixed $\gamma$ is when $k=\lfloor\frac{m}{2}-1\rfloor$ leading to a value for the displacement field of
$u=\gamma\frac{2(m-1)}{\sqrt{(m-3)(m-1)}}.$
For $m$ large, this is asymptotically $u\approx 2\gamma$, which is significantly larger than the values displayed in current experiments [17], which are constrained to around $u\approx 0.221$ eV by a sufficient spin-orbit coupling (SOC) ensuring an appropriate spin polarization. The largest critical value of $u$ is given by (13) with $\delta_{k}=1$ when $m\geq 4$ is even and $\delta_{k}=\frac{3}{2}$ when $m\geq 5$ is odd so that the largest transition value of $u$ is proportional to $m\gamma$.
The main result of this paper is the following:
Theorem 2.3 (QAH phases)
@@
${\rm BDI}\,(\gamma,u)={\rm BDI}_{j}:=\frac{m^{2}}{2}-2\delta_{j}(\delta_{j}+1)\quad\mbox{ when }\quad\delta_{j}<\delta(\gamma,u)<\delta_{j+1}$ (14)
for $0\leq j\leq\lfloor\frac{m}{2}-1\rfloor$ and $\delta_{j}$ given by (12).
@@
On the other hand, when $u$ is so large that $\delta_{0}\leq\frac{1}{2}<\delta(\gamma,u)<\delta_{1}$, then the QAHE is maximal for a given value of $m$ and given for $j=0$ by $\frac{1}{2}(m^{2}-3)$ when $m$ is odd and by $\frac{1}{2}m^{2}$ when $m$ is even. The result for $m=2n+1$ corresponds to the $n$-replica model analyzed in [7].
For concreteness, we observe the following possible values of the BDI for $2\leq m\leq 11$:
m=  2 3 4 5 6 7 8 9 10 11
8 11 14 17 20 23 26 29
18 23 28 33 38 43
32 39 46 53
50 59
Figure 1: Possible values for the BDI of the QAHE with $2\leq m\leq 11$ layers.
The values of the invariants in (14) were computed for a valley index $\tau=+1$ and a Bernal stacking $B=\gamma A$. Using (8) and the fact that $U(0)$ in (19) below is replaced by $U(0)^{\prime}$ when $U(\theta)$ is replaced by $U^{*}(\theta)$, we find that the above invariants satisfy
${\rm BDI}_{j}[1,\gamma A]=-{\rm BDI}_{j}[-1,\gamma A]=-{\rm BDI}_{j}[1,\gamma A^{*}]={\rm BDI}_{j}[-1,=\gamma A^{*}],$
where we use the notation ${\rm BDI}_{j}={\rm BDI}_{j}[\tau,B]$. For a given expression for $B$, we thus obtain that ${\rm BDI}_{j}[\tau=1]+{\rm BDI}_{j}[\tau=-1]=0$, implying that the topological invariant accounting for both valleys $\tau=\pm 1$ is always trivial.
@@
Following (10), the derivation is based on computing the integral of Berry curvature associated to $\hat{H}^{h}(\mathbf{k})$ for $h=N/S$. Let $\hat{H}({\mathbf{k}})=\sum_{j}E_{j}(k)\Pi_{j}({\mathbf{k}})$ denote one of them. Following Theorem 2.2, the spectrum of this Hamiltonian is invariant by rotation with $E_{j}(k)$ simple eigenvalues, i.e., associated to $\Pi_{j}({\mathbf{k}})$ a rank-one projector, and so that $E_{2m+1-j}(k)=-E_{j}(k)$ for $1\leq j\leq 2m$.
@@
Computing the integrals appearing in (10) analytically drastically simplifies in the presence of rotational symmetry. The Chern number of the projectors $\Pi^{h}({\mathbf{k}})$ is additive and hence may be written as the sum over branches $E_{j}(k)<0$ of the rank-one projectors $\Pi_{j}({\mathbf{k}})$ [9].
@@
Assume $\Pi({\mathbf{k}})=\psi({\mathbf{k}})\psi^{*}({\mathbf{k}})$ is such a rank-one projector dropping the index $j$ for simplification. We then verify that
${\rm tr}\Pi d\Pi\wedge d\Pi=dA,\qquad A({\mathbf{k}})=(\psi({\mathbf{k}}),d\psi({\mathbf{k}}))$
@@
where $A$ is a one-form-$i\mathbb{R}$-valued (Berry) connection assuming $d\psi({\mathbf{k}})$ continuously defined, which can always be achieved since $\mathbb{R}^{2}$ is contractible. As an application of the Stokes theorem, we thus observe [9, 28, 18] that
$\mathcal{C}[\Pi]=\frac{i}{2\pi}\Big{(}\oint_{|{\mathbf{k}}|\rightarrow\infty}A({\mathbf{k}})-\oint_{|{\mathbf{k}}|\rightarrow 0}A({\mathbf{k}})\Big{)}.$ (15)
@@
When $A({\mathbf{k}})$ is continuously defined at ${\mathbf{k}}=0$, then the above integral over circles with vanishingly small radii converges to 0. However, writing ${\mathbf{k}}=|{\mathbf{k}}|e^{i\theta}$ for $\theta\in[0,2\pi)$, we observe that the above formula remains valid if $\psi({\mathbf{k}})$ is (globally) gauge-transformed to, e.g., $e^{im\theta}\psi({\mathbf{k}})$. This flexibility proves convenient in practice.
Assume now that the Hamiltonian family is isotropic. Let ${\mathbf{k}}=ke^{i\theta}$ for $e^{i\theta}\in\mathbb{S}^{1}$ the unit circle. We know that
$H({\mathbf{k}})=U(\theta)H(0)U^{*}(\theta)$
@@
for $U(\theta)$ a family of $\mathbb{C}^{n}$- unitary transformations. This implies that the branches of spectrum $\lambda_{j}({\mathbf{k}})=\lambda_{j}(k)$ are independent of $\theta$ and we may choose the eigenvectors as $\psi_{j}(\theta)=U(\theta)\psi_{j}(0)$.
@@
Isotropy, or invariance by rotation, implies that $U^{*}(\theta)dU(\theta)=U^{\prime}(0)d\theta$ is independent of $\theta\in[0,2\pi)$. In this setting, we thus obtain that
$\mathcal{C}[\Pi]=i\big{[}\lim_{k\to\infty}(\psi(k),U^{\prime}(0)\psi(k))-\lim_{k\to 0}(\psi(k),U^{\prime}(0)\psi(k))\big{]}.$ (16)
@@
In other words, all we need to compute is $\psi({\mathbf{k}})$ for ${\mathbf{k}}=(k,0)$ with $k=0$ and $k=\infty$, which may be obtained analytically. Explicit expressions for the projectors as $k\to\infty$ are also necessary, for instance in order to verify gluing conditions recalled below in (25).
Eigenvectors of $H_{B}^{N/S}$ as $k\to\infty$.
@@
Suppose that $\psi=(a_{1},b_{1},a_{2},b_{2},...,a_{m},b_{m})^{t}$ is an eigenvector of $H_{B}$ with eigenvalue $\lambda$; $H_{B}\psi=\lambda\psi$. First consider the $2j$ and 2j12j-1 lines of the eigenvalue equation:
@@
From Theorem 2.2 we know that there must be $m$ positive eigenvalues for $k\neq 0$, and furthermore since $\hat{H}_{B}^{h}$ is Hermitian $\{\psi_{j}^{+}\}_{j=1}^{m}$ is an orthonormal set. Now clearly from the above expression for $\psi_{j}^{+}$, $\{\phi_{j}\}_{j=1}^{m} = span$$\{\psi_{j}^{+}\}_{j=1}^{m}$. Therefore setting:
$\Pi_{j}^{h}=\psi_{j}^{h}\otimes\psi_{j}^{h},\qquad\tilde{\Pi}_{j}^{h}=\phi_{j}^{h}\otimes\phi_{j}^{h}$
@@
$I-P^{h}_{m}=\sum_{j\geq m+1}\Pi_{j}^{h}=\sum_{j\geq m+1}\tilde{\Pi}_{j}^{h}.$
Since ${\mathfrak{C}}_{l}=\mathcal{W}_{l}^{S}-\mathcal{W}_{l}^{N}=\mathcal{C}[P^{S}_{l}]-\mathcal{C}[P_{l}^{N}]$, the BDI is independent of the choice of basis in which we choose to express $P_{l}^{h}$. In other words, since we have found that $\text{Ran}(P_{m}^{h})$ has an orthonormal basis $\{\phi_{j}\}_{j=1}^{m}$ we are free to use $\phi_{j}$ in place of $\psi_{j}^{+}$ in (16) provided we are summing over $j=\{1,...,m\}$ to obtain ${\mathfrak{C}}_{m}$.
@@
$\lambda_{j}^{\pm}\approx u_{j}\pm\alpha_{j}|k|,\qquad\psi_{j}^{\pm}=\frac{1}{\sqrt{2}}(\hat{e}_{2j-1}\pm\hat{e}_{2j}).$
It remains to justify the replacement of each $k\sigma_{1}$ by $0,1]\ni t\mapsto((1-t)+t\alpha_{n})k\sigma_{1}$, which is continuous in $t$. For each value of $t$, the corresponding interface Hamiltonian satisfies [H1] in [3, 26] so that the BDI is given by the Fedosov-Hörmander formula [3, Theorem 4.13], an integer-valued topological winding number that is clearly independent of $t\in[0,1]$ (see, [4] for more detail). These expressions allow us verify gluing conditions (25) without finding eigenvectors which in fact rely non-trivially on $(u,\gamma)$, although solving (18) explicitly is also possible to verify (25), as done in [6] for the 2-layer case.
Regardless of the approach taken, we have shown that the set of positive (or negative) eigenvectors as $k\to\infty$ does not depend on the sign of $u$, and therefore the term $\lim_{k\to\infty}(\psi(k),U^{\prime}(0)\psi(k))$ in (16) does not contribute to the BDI for any parameter values after summing over $\sum_{j\leq m}\mathcal{C}[\Pi_{j}^{h}]$.
Eigenvectors of $H_{B}^{N/S}$ as $k\to 0$.
For $k=0$ we note the following eigenvectors which we are able to be computed explicitly:
$E_{1}=u_{1},\;\;\psi_{1}(k=0)=(1,0,0,...,0)^{t}$
$E_{m}=u_{m},\;\;\psi_{m}(k=0)=(0,0,0,...,0,1)^{t}$
$E_{j}^{\pm}=\frac{u}{m-1}\Big{[}\big{(}j-\frac{m}{2}\big{)}\pm\delta(\gamma,u)\Big{]},\;\;\;\psi_{j}^{+}(k=0)=c\hat{e}_{2j}+s\hat{e}_{2j+1},\;\;\;\psi_{j}^{-}(k=0)=c\hat{e}_{2j+1}-s\hat{e}_{2j},$
for $2\leq j\leq m-1$ and $c^{2}+s^{2}=1$. We obtain from (5):

### v2

@@ 2 Effective model and classification
for $0\leq k\leq\lfloor\frac{m}{2}\rfloor$. For cases we consider below where $k<0$ we define $\delta_{k}=\delta_{-k}$. Setting $j=k+\lceil\frac{m}{2}\rceil$ and rearranging (11) in terms of $\delta_{k}$ we deduce that phase transitions happen at critical points of potential $u$, assuming $\gamma$ remains constant:
$u_{\pm k}=\pm\gamma\frac{m-1}{\sqrt{\delta_{k}^{2}-\frac{1}{4}}}$ (13)
for $1\leq k\leq\lfloor\frac{m}{2}-1\rfloor$. Noting that $u=0$ is also a critical value (where in fact all eigenvalues coincide at 0 when $\mathbf{k}=0$)[27], there are therefore $2\lfloor\frac{m}{2}\rfloor$ distinct bulk topological phases which we label by $j\in\{\pm 1,\pm 2,...,\pm\lfloor\frac{m}{2}\rfloor\}$ defined by:
$\hat{H}_{B}(\mathbf{k},\gamma,u)\in\text{sgn}(u)j\quad\text{iff}\quad\delta_{j-1}<\delta(\gamma,u)<\delta_{j};\;\;1\leq j\leq\lfloor\frac{m}{2}\rfloor-1$ (14)
and $\hat{H}_{B}(\mathbf{k},\gamma,u)\in\text{sgn}(u)\lfloor\frac{m}{2}\rfloor$ iff $\delta(\gamma,u)>\delta_{\lfloor\frac{m}{2}\rfloor-1}$ (phase $\lfloor\frac{m}{2}\rfloor$ has no upper bound in $\delta(\gamma,u))$.
We thus observe the existence of only two phases when $m=2$ or $m=3$. Assuming $m\geq 4$, the smallest value of $u$ at which we may observe a transition for a fixed $\gamma$ is when $k=\lfloor\frac{m}{2}-1\rfloor$ leading to first critical value for the displacement field of
$u_{\lfloor\frac{m}{2}-1\rfloor}=\gamma\frac{2(m-1)}{\sqrt{(m-3)(m-1)}}.$
For $m$ large, this is asymptotically $u\approx 2\gamma$, which is significantly larger than the values displayed in current experiments [21, 11], which are constrained to around $u\approx 0.221$ eV and $u\approx 0.209$ eV respectively by a sufficient spin-orbit coupling (SOC) ensuring an appropriate spin polarization (recall that $\gamma$ is estimated to be 0.435 eV [21]). The largest critical value of $u$ is given by (13) with $\delta_{k}=1$ when $m\geq 4$ is even and $\delta_{k}=\frac{3}{2}$ when $m\geq 5$ is odd so that the largest transition value of $u$ is proportional to $m\gamma$.
We are now ready to state the main result of the paper, the proof of which is delayed to the next section:
Theorem 2.3 (BDI for QAH phases)
@@
${\rm BDI}_{kj}=\begin{cases}\text{sgn}(j)\left(\delta_{k}(\delta_{k}-1)-\delta_{j}(\delta_{j}-1)\right)&\text{sgn}(j)=\text{sgn}(k)\\ \text{sgn}(j)\left(\frac{m^{2}}{2}-(\delta_{j}(\delta_{j}-1)+\delta_{k}(\delta_{k}-1))\right)&\text{sgn}(j)=-\text{sgn}(k).\end{cases}$ (15)
for $j,k\in\{\pm 1,\pm 2,...,\pm\lfloor\frac{m}{2}\rfloor\}$ and $\delta_{j}$ given by (12).
@@
Keeping in mind that bulk phases with gate potential u-u and $u$ are energetically equivalent we consider the transition between phases k-k and $j$ to be equivalent to the transition between j-j and $k$ (note the equivalence of (15) for these two these transitions). Also note that ${\rm BDI}_{kj}=-{\rm BDI}_{jk}$. Disregarding these degeneracies we find $\lfloor\frac{m}{2}\rfloor^{2}$ distinct transitions between QAH phases. For concreteness, Figure 1 enumerates all possible absolute values of BDI’s according to (15) for a range of values $2\leq m\leq 9$.
m = 2
3
2 4 6 8
3 5 8 11
2 4 6∗† 10 12 14 16 18
3 5 7 8 12 15 17 20 23
2 4 6* 8 10 12 14 18 20∗† 24 26 28 30 32
3 5 7 8 9 12 15 16 21 23 24 28 31 33 36 39
Figure 1: All possible values of the BDI for the QAHE with $2\leq m\leq 9$. The symmetric case in which $u^{N}=-u^{S}>0$ is denoted by ; see Figure 4. Bold numbers in each row reflect the number of layers m in each row (which is always a BDI). An * reflects that there is a degeneracy of two topological transitions with the same BDI which are not energetically equivalent.
The values of the invariants in (15) were computed for a valley index $\tau=+1$ and ABC stacking $B=\gamma A$. Using (8) and the fact that $U^{\prime}(0)$ in (20) below is replaced by $U(0)^{\prime}$ when $U(\theta)$ is replaced by $U^{*}(\theta)$, we find that the above invariants satisfy
${\rm BDI}_{kj}[1,\gamma A]=-{\rm BDI}_{kj}[-1,\gamma A]=-{\rm BDI}_{kj}[1,\gamma A^{*}]={\rm BDI}_{kj}[-1,\gamma A^{*}],$
where we use the notation ${\rm BDI}_{kj}={\rm BDI}_{kj}[\tau,B]$. For a given expression for $B$, we thus obtain that ${\rm BDI}_{j}[\tau=1]+{\rm BDI}_{j}[\tau=-1]=0$, implying that the topological invariant accounting for both valleys $\tau=\pm 1$ is always trivial in the model we consider here.
@@
In this section we prove Theorem 2.3. Following (10), the derivation is based on computing the integral of Berry curvature associated to $\hat{H}^{h}(\mathbf{k})$ for $h=N/S$. Let $\hat{H}({\mathbf{k}})=\sum_{j}E_{j}(k)\Pi_{j}({\mathbf{k}})$ denote one of them. Following Theorem 2.2, the spectrum of this Hamiltonian is invariant by rotation with $E_{j}(k)$ simple eigenvalues, i.e., associated to $\Pi_{j}({\mathbf{k}})$ a rank-one projector, and so that $E_{2m+1-j}(k)=-E_{j}(k)$ for $1\leq j\leq 2m$. In this section we take the view that $\mathbf{k}\in\mathbb{R}^{2}$ to be consistent with (10) and the supporting theory, and identify $(k,\theta)$ as the polar coordinates of $\mathbf{k}\in\mathbb{R}^{2}$.
@@
Computing the integrals appearing in (10) analytically drastically simplifies in the presence of rotational symmetry. The Chern number of the projectors $\Pi^{h}(k,\theta)$ (recall (9)) is additive and hence may be written as the sum over branches $E_{j}(k)<0$ of the rank-one projectors $\Pi_{j}(k,\theta)$ onto the eigenspace associated to $E_{j}(k)$[9].
@@
Assume $\Pi_{j}(k,\theta)=|\psi_{j}(k,\theta)\rangle\langle\psi_{j}(k,\theta)|$ is such a rank-one projector. We then verify that
${\rm tr}\Pi_{j}d\Pi_{j}\wedge d\Pi_{j}=dA_{j},\qquad A_{j}(k,\theta)=(\psi_{j}(k,\theta),d\psi_{j}(k,\theta))$
@@
where $A$ is a one-form-$i\mathbb{R}$-valued (Berry) connection assuming $d\psi_{j}(k,\theta)$ continuously defined, which can always be achieved since $\mathbb{R}^{2}$ is contractible. As an application of the Stokes theorem, we thus observe [9, 32, 22] that
$\mathcal{C}[\Pi_{j}]:=\frac{i}{2\pi}\int_{\mathbb{R}^{2}}dA_{j}(k,\theta)=\frac{i}{2\pi}\Big(\oint_{k\rightarrow\infty}A_{j}(k,\theta)-\oint_{k\rightarrow 0}A_{j}(k,\theta)\Big).$ (16)
@@
When $A_{j}(k,\theta)$ is continuously defined at ${\mathbf{k}}=0$, then the above integral over circles with vanishingly small radii converges to 0. However, since $\psi_{j}(k,\theta)$ is an eigenvector of $H(k,\theta)$ we observe that the above formula remains valid if $\psi_{j}(k,\theta)$ is (globally) gauge-transformed to, e.g., $e^{im\theta}\psi_{j}(k,\theta)$. This flexibility proves convenient in practice.
We can further simplify (16) by taking advantage of the isotropy of $\hat{H}(k,\theta)$. We know that
$\hat{H}(k,\theta)=U(\theta)\hat{H}(k,0)U^{*}(\theta)$
@@
for $U(\theta)$ a family of $\mathbb{C}^{n}$- unitary transformations. This implies that the branches of spectrum $\lambda_{j}(k,\theta)=\lambda_{j}(k)$ are independent of $\theta$ and we may choose the eigenvectors as $\psi_{j}(k,\theta)=U(\theta)\psi_{j}(k,0)$. Thus we obtain:
$A_{j}(k,\theta)=\Big(U(\theta)\psi_{j}(k,0),d\big(U(\theta)\psi_{j}(k,0)\big)\Big)\qquad\qquad\qquad\qquad\qquad\qquad\qquad$
$\qquad=\Big(\psi_{j}(k,0),U^{*}(\theta)dU(\theta)\psi_{j}(k,0)\Big)+\big(\psi_{j}(k,0),\frac{d}{dk}\psi_{j}(k,0)\big)dk.$
@@
Isotropy, or invariance by rotation, implies that $U^{*}(\theta)dU(\theta)=U^{\prime}(0)d\theta$ is independent of $\theta\in[0,2\pi)$ (which can also be verified directly from (5)). In this setting, we thus obtain, since the integrals in (16) are taken on constant $k$ curves, that
$\mathcal{C}[\Pi_{j}]=i\big[\lim_{k\to\infty}(\psi_{j}(k,0),U^{\prime}(0)\psi_{j}(k,0))-\lim_{k\to 0}(\psi_{j}(k,0),U^{\prime}(0)\psi_{j}(k,0))\big].$ (17)
@@
In other words, all we need to compute is $\psi_{j}(k,0)$ with $k\to 0$ and $k\to\infty$, which may be obtained analytically, to obtain ${\rm BDI}(u^{S},u^{N},\gamma)={\mathfrak{C}}[\Pi^{S},\Pi^{N}]=\mathcal{C}[\Pi^{S}]-\mathcal{C}[\Pi^{N}]=\sum_{j\leq m}\left(\mathcal{C}[\Pi_{j}^{S}]-\mathcal{C}[\Pi_{j}^{N}]\right)$ by the (non-trivial) additivity of Chern numbers. Explicit expressions for the projectors as $k\to\infty$ are also necessary in order to verify gluing conditions recalled below in (27).
Eigenvectors of $\hat{H}_{B}^{N/S}$ as $k\to\infty$.
@@
Suppose that $\psi(k,0)=(a_{1},b_{1},a_{2},b_{2},...,a_{m},b_{m})^{t}$ is an eigenvector of $\hat{H}$ with eigenvalue $\lambda$; $\hat{H}(k,0)\psi=\lambda\psi(k,0)$. First consider the $2j$ and 2j12j-1 lines of the eigenvalue equation:
@@
From Theorem 2.2 we know that there must be $m$ positive eigenvalues for $k\neq 0$, and furthermore since $\hat{H}$ is Hermitian $\{\psi_{j}^{+}\}_{j=1}^{m}$ is an orthonormal set. Now clearly from the above expression for $\psi_{j}^{+}$, $\{\phi_{j}\}_{j=1}^{m} = span$$\{\psi_{j}^{+}\}_{j=1}^{m}$. Therefore setting:
$\Pi_{j}^{h}=|\psi_{j}^{h}\rangle\langle\psi_{j}^{h}|,\qquad\tilde{\Pi}_{j}^{h}=|\phi_{j}^{h}\rangle\langle\phi_{j}^{h}|$
@@
$I-\Pi^{h}=\sum_{j\geq m+1}\Pi_{j}^{h}=\sum_{j\geq m+1}\tilde{\Pi}_{j}^{h}.$
Since ${\mathfrak{C}}[\Pi^{S},\Pi^{N}]=\mathcal{C}[\Pi^{S}]-\mathcal{C}[\Pi^{N}]$, the BDI is independent of the choice of basis in which we choose to express $\Pi^{h}$. In other words, since we have found that $\text{Ran}(\Pi^{h})$ has an orthonormal basis $\{\phi_{j}\}_{j=1}^{m}$ we are free to use $\phi_{j}$ in place of $\psi_{j}^{+}$ in (17) provided we are summing over $j=\{1,...,m\}$ to obtain $\mathcal{C}[\Pi^{S}],\mathcal{C}[\Pi^{N}]$.
@@
$\lambda_{j}^{\pm}\approx u_{j}\pm\alpha_{j}k,\qquad\psi_{j}^{\pm}=\frac{1}{\sqrt{2}}(\hat{e}_{2j-1}\pm\hat{e}_{2j}).$
It remains to justify the replacement of each $k\sigma_{1}$ by $0,1]\ni t\mapsto((1-t)+t\alpha_{n})k\sigma_{1}$, which is continuous in $t$. For each value of $t$, the corresponding interface Hamiltonian satisfies [H1] in [6, 30] so that the BDI is given by the Fedosov-Hörmander formula [6] (Theorem 4.13), an integer-valued topological winding number that is clearly independent of $t\in[0,1]$[7]. These expressions allow us verify gluing conditions (27) without finding eigenvectors which in fact rely non-trivially on $(u,\gamma)$, although solving (19) explicitly is also possible to verify (27), as done in [2] for the 2-layer case.
Regardless of the approach taken, we have shown that the subspace spanned by the positive (or negative) eigenvectors as $k\to\infty$ does not depend on the parameters $(\gamma,u)$, and therefore the term $\lim_{k\to\infty}(\psi(k,0),U^{\prime}(0)\psi(k,0))$ in (17) after summing over $\sum_{j\geq m+1}\mathcal{C}[\Pi_{j}^{h}]$ or $\sum_{j\leq m}\mathcal{C}[\Pi_{j}^{h}]$ is identical for any phase (for the latter simply repeat the preceding arguments with $\lambda<0$). In particular denoting $\sum_{j\leq m}\mathcal{C}[\Pi_{j}^{h}]=c_{\infty}$ we may add a global gauge transformation $e^{-ic_{\infty}\theta}$ to the eigenvectors $\psi_{j}$ so that the contribution of the $k\to\infty$ terms to $\mathcal{C}[\Pi^{h}]$ is in fact 0 for all phases.
Eigenvectors of $\hat{H}_{B}^{N/S}$ as $k\to 0$.
For $k=0$ we note the following eigenvectors:
$E_{0}=u_{1}=-\frac{u}{2},\;\;\psi_{0}(0,0)=(1,0,0,...,0)^{t}$
$E_{m}=u_{m}=\frac{u}{2},\;\;\psi_{m}(0,0)=(0,0,0,...,0,1)^{t}$
$E_{n}^{\pm}=\frac{u}{m-1}\Big[\big(n-\frac{m}{2}\big)\pm\delta(\gamma,u)\Big],\;\;\;\psi_{n}^{+}(0,0)=c\hat{e}_{2n}+s\hat{e}_{2n+1},\;\;\;\psi_{n}^{-}(0,0)=c\hat{e}_{2n+1}-s\hat{e}_{2n},$
for $1\leq n\leq m-1$ and $c^{2}+s^{2}=1$. We obtain from (5):